# Gesture Recognition from body-Worn RFID under Missing Data

### v1

@@ II Related Work
Vectors and matrices are denoted by lowercase and uppercase bold letters, respectively. The $i$-th element of vector $\mathbf{a}$ is represented by $\mathbf{a}]_{i}$. For a matrix $\mathbf{A}$, $\mathbf{A}(:,i:j)$ selects all elements in all rows and columns $i$ through $j$ (inclusive), while $\mathbf{A}(i:j,:)$ selects all elements in rows $i$ through $j$ (inclusive) and all columns. The elements in the $\ell$-th row and $\ell$-th column are denoted by $\mathbf{A}(\ell,:)$ and $\mathbf{A}(:,\ell)$, respectively. For a vector $\mathbf{a}$, we use the notation $\mathbf{a}_{i:j}$ to denote the subsequence consisting of the elements from position $i$ to position $j$ (inclusive). The transpose operator is $(\cdot)^{\top}$. The $L_{2}$ norm (Euclidean norm for vectors, Frobenius norm for matrices) is denoted by $\|\cdot\|_{2}$. The cardinality of a set or the length of a vector is denoted by $|\cdot|$. The $L_{0}$ norm, $\|\mathbf{a}\|_{0}$, denotes the number of non-zero elements in $\mathbf{a}$. A column vector of length $Q$ with all elements equal to one is represented by $\mathbf{1}_{Q}$, and a column vector or matrix of size $Q$ with all elements equal to zero is represented by $\mathbf{0}_{Q}$. Assignment, where the value of $b$ replaces the value of $a$, is denoted by $a\leftarrow b$. In this paper, for two vectors $\mathbf{a}$ and $\mathbf{b}$, we define $\mathbf{a}\setminus\mathbf{b}$ as a vector containing the elements of $\mathbf{a}$ that are not present in $\mathbf{b}$, preserving their original order from $\mathbf{a}$. Note that we assume $\mathbf{b}$ is a sub-sequence of $\mathbf{a}$ for this operation. For example, if $\mathbf{a}=[1,2,3,4]$ and $\mathbf{b}=[1,4]$, then $\mathbf{a}\setminus\mathbf{b}=[2,3]$. The function $\operatorname{argsort}(\mathbf{a})$ returns the indices that would sort the vector $\mathbf{a}$ in ascending order. The complement of a set $a$ is denoted by $a^{c}$, and $\emptyset$ denotes the empty set. For two sets $A$ and $B$, the notation $A\subseteq B$ means that $A$ is a subset of $B$.
@@ III Overview of the measurement System
We utilize Alien AZ 9662 passive RFID tags as well as an Impinj Speedway R420 RFID reader and a circularly polarized Vulcan RFID PAR90209H antenna with a circular polarization gain of $9$ dBiC as well as elevation and azimuth beamwidth of $70^{\circ}$. The reader is controlled by a laptop running the Impinj ItemTest software. The RFID system operates in the 865MHz range at $30$ dBm.
The integrated circuit (IC) of the tag modulates its response to the continuous wave signal emitted by the reader by altering the load on its antenna to either backscatter or absorb the reader’s signal, thereby producing reflection coefficient $\Gamma(t)\in\{0,1\}$. The reflection coefficient is $\Gamma=\frac{Z_{\text{Ant}}-Z_{\text{Load}}(t)}{Z_{\text{Ant}}+Z_{\text{Load}}(t)}$, where $Z_{\text{Ant}}$ and $Z_{\text{Load}}(t)$ denote the antenna impedance (around $50\Omega$) and the load impedance controlled by the tag’s IC. $Z_{\text{Load}}=Z_{\text{Ant}}$ will result in $\Gamma=0$ (no signal reflected) while $Z_{\text{Load}}=0$ ($\Gamma=1$) causes the signal of the reader to be reflected. [^1] In practice, due to imperfect load matching, the maximum value of the $\Gamma$ is close to $1$ and the minimum value is close to zero.
@@ IV-B Data Imputation
In the first case, the $i$-th reference gesture data contains only one null dataframe for the EPC with index $p$. We first calculate all MED values between the $i$-th reference gesture data and the other gesture data within the same class, i.e., $\kappa_{i,m}\quad\forall m\neq i$ given in (9). We then define the matrix $\bm{K}$, where each row contains $\kappa_{i,m}$ in the first column and the corresponding index $m$ in the second column. We identify the gesture data indices corresponding to the $\nu$ smallest values in the first column of $\bm{K}$ (i.e., $\bm{K}(:,1)$) by sorting it in ascending order. Let $v_{1}\leq v_{2}\leq\dots\leq v_{\nu}$ be these $\nu$ smallest MED values, and let $\bm{\zeta}=[\zeta_{1},\zeta_{2},\dots,\zeta_{\nu}]^{\top}$ be the vector of the corresponding gesture data indices, such that, for example, $\kappa_{i,\zeta_{1}}=v_{1}$.
Then, we define the set $\Omega\triangleq\{\forall\zeta_{j}\mid\tilde{\mathbf{D}}^{p}_{\zeta_{j}}(:,3:4)\neq\mathbf{0},j=1,2,\dots,\nu\}$, where $\tilde{\mathbf{D}}^{p}_{\zeta_{j}}(:,3:4)\neq\mathbf{0}$ indicates that the $\zeta_{j}$-th gesture in $\Omega$ contains a sparse dataframe for the EPC with index $p$. Finally, the missing value $\tilde{\mathbf{D}}^{p}_{i}(:,3:4)$ is imputed by averaging the RSS and phase values from the set $\Omega$:
@@
0:$i$-th raw gesture data $\mathbf{D}_{i}\in\mathbb{R}^{T\times 4}$, $N$, and $l_{\text{rs}}$ Extract $i$-th sorted gesture data, $\tilde{\mathbf{D}}_{i}$, and dataframes $\tilde{\mathbf{D}}^{n}_{i}$ for all $n\in\{1,2,\dots,N\}$
@@ V Attention-Based Graph Learning
To facilitate model training, the input data is structured into a four-dimensional tensor $\bm{\mathcal{T}}\in\mathbb{R}^{B\times T\times N\times D}$ where $B$ denotes the total number of gesture samples, $T$ is the number of timestamps per EPC in each gesture sample, $N$ represents the number of EPCs, and $D$ is the feature dimension of each EPC (i.e., RSS and phase). In this work, the tensor shape used is ($T=30$, $N=8$, $D=2$), corresponding to 30 timestamps, 8 EPCs, and two-dimensional feature vectors per node.

### v2

@@ II Related Work
Vectors and matrices are denoted by lowercase and uppercase bold letters, respectively. The $i$-th element of vector $\mathbf{a}$ is represented by $\mathbf{a}]_{i}$. For a matrix $\mathbf{A}$, $\mathbf{A}(:,i:j)$ selects all elements in all rows and columns $i$ through $j$ (inclusive), while $\mathbf{A}(i:j,:)$ selects all elements in rows $i$ through $j$ (inclusive) and all columns. The elements in the $\ell$-th row and $\ell$-th column are denoted by $\mathbf{A}(\ell,:)$ and $\mathbf{A}(:,\ell)$, respectively. For a vector $\mathbf{a}$, we use the notation $\mathbf{a}_{i:j}$ to denote the subsequence consisting of the elements from position $i$ to position $j$ (inclusive). The transpose operator is $(\cdot)^{\top}$. The $L_{2}$ norm (Euclidean norm for vectors, Frobenius norm for matrices) is denoted by $\|\cdot\|_{2}$. The cardinality of a set or the length of a vector is denoted by $|\cdot|$. The $L_{0}$ norm, $\|\mathbf{a}\|_{0}$, denotes the number of non-zero elements in $\mathbf{a}$. A column vector of length $Q$ with all elements equal to one is represented by $\mathbf{1}_{Q}$, and a column vector or matrix of size $Q$ with all elements equal to zero is represented by $\mathbf{0}_{Q}$. Assignment, where the value of $b$ replaces the value of $a$, is denoted by $a\leftarrow b$. In this paper, for two vectors $\mathbf{a}$ and $\mathbf{b}$, we define $\mathbf{a}\setminus\mathbf{b}$ as a vector containing the elements of $\mathbf{a}$ that are not present in $\mathbf{b}$, preserving their original order from $\mathbf{a}$. Note that we assume $\mathbf{b}$ is a sub-sequence of $\mathbf{a}$ for this operation. For example, if $\mathbf{a}=[1,2,3,4]$ and $\mathbf{b}=[1,4]$, then $\mathbf{a}\setminus\mathbf{b}=[2,3]$. The function $\operatorname{argsort}(\mathbf{a})$ returns the indices that would sort the vector $\mathbf{a}$ in ascending order. The complement of a set $a$ is denoted by $a^{c}$, and $\emptyset$ denotes the empty set. For two sets AA and $B$, the notation $A\subseteq B$ means that AA is a subset of $B$.
@@ III Overview of the measurement System
We utilize Alien AZ 9662 passive RFID tags as well as an Impinj Speedway R420 RFID reader and a circularly polarized Vulcan RFID PAR90209H antenna with a circular polarization gain of 99 dBiC as well as elevation and azimuth beamwidth of $70^{\circ}$. The reader is controlled by a laptop running the Impinj ItemTest software. The RFID system operates in the 865MHz range at $30$ dBm.
The integrated circuit (IC) of the tag modulates its response to the continuous wave signal emitted by the reader by altering the load on its antenna to either backscatter or absorb the reader’s signal, thereby producing reflection coefficient $\Gamma(t)\in\{0,1\}$. The reflection coefficient is $\Gamma=\frac{Z_{\text{Ant}}-Z_{\text{Load}}(t)}{Z_{\text{Ant}}+Z_{\text{Load}}(t)}$, where $Z_{\text{Ant}}$ and $Z_{\text{Load}}(t)$ denote the antenna impedance (around $50\Omega$) and the load impedance controlled by the tag’s IC. $Z_{\text{Load}}=Z_{\text{Ant}}$ will result in $\Gamma=0$ (no signal reflected) while $Z_{\text{Load}}=0$ ($\Gamma=1$) causes the signal of the reader to be reflected. [^1] In practice, due to imperfect load matching, the maximum value of the $\Gamma$ is close to 11 and the minimum value is close to zero.
@@ IV-B Data Imputation
In the first case, the $i$-th reference gesture data contains only one null dataframe for the EPC with index pp. We first calculate all MED values between the $i$-th reference gesture data and the other gesture data within the same class, i.e., $\kappa_{i,m}\quad\forall m\neq i$ given in (9). We then define the matrix $\bm{K}$, where each row contains $\kappa_{i,m}$ in the first column and the corresponding index $m$ in the second column. We identify the gesture data indices corresponding to the $\nu$ smallest values in the first column of $\bm{K}$ (i.e., $\bm{K}(:,1)$) by sorting it in ascending order. Let $v_{1}\leq v_{2}\leq\dots\leq v_{\nu}$ be these $\nu$ smallest MED values, and let $\bm{\zeta}=[\zeta_{1},\zeta_{2},\dots,\zeta_{\nu}]^{\top}$ be the vector of the corresponding gesture data indices, such that, for example, $\kappa_{i,\zeta_{1}}=v_{1}$.
Then, we define the set $\Omega\triangleq\{\forall\zeta_{j}\mid\tilde{\mathbf{D}}^{p}_{\zeta_{j}}(:,3:4)\neq\mathbf{0},j=1,2,\dots,\nu\}$, where $\tilde{\mathbf{D}}^{p}_{\zeta_{j}}(:,3:4)\neq\mathbf{0}$ indicates that the $\zeta_{j}$-th gesture in $\Omega$ contains a sparse dataframe for the EPC with index pp. Finally, the missing value $\tilde{\mathbf{D}}^{p}_{i}(:,3:4)$ is imputed by averaging the RSS and phase values from the set $\Omega$:
@@
0:$i$-th raw gesture data $\mathbf{D}_{i}\in\mathbb{R}^{T\times 4}$, NN, and $l_{\text{rs}}$ Extract $i$-th sorted gesture data, $\tilde{\mathbf{D}}_{i}$, and dataframes $\tilde{\mathbf{D}}^{n}_{i}$ for all $n\in\{1,2,\dots,N\}$
@@ V Attention-Based Graph Learning
To facilitate model training, the input data is structured into a four-dimensional tensor $\bm{\mathcal{T}}\in\mathbb{R}^{B\times T\times N\times D}$ where $B$ denotes the total number of gesture samples, $T$ is the number of timestamps per EPC in each gesture sample, NN represents the number of EPCs, and $D$ is the feature dimension of each EPC (i.e., RSS and phase). In this work, the tensor shape used is ($T=30$, $N=8$, $D=2$), corresponding to 30 timestamps, 8 EPCs, and two-dimensional feature vectors per node.